Determine if each is a function: 20. y13 = x + 5- 21. x4 = y +45 22. x19.+ 4y18 = 16 23. xº + y2 = 10
Answer #1

Note that a relation fixy is a function . if for every x in x there is unique Yin Y such that fra)=y. or for giren egn if we22) gely + 4y18=16 = 4y8 = 16-gely * . yl8 = 16-os y29 = 16-219 9 ye = ( 15-xely ys 4= + (16-809 ) :. Given equation is not a